40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12526     Entry time: Fri Sep 30 19:53:07 2016     In reply to: 12122     Reply to this: 12532
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: X end IR pickoff fiber coupled 

[johannes, gautam]

Today we re-installed the fiber coupler on the X-endtable to couple some of the PSL light into a fiber that runs to the PSL table, where it is combined with a similar PSL pickoff to make an IR beat between the EX AUX laser and the PSL. The main motivation behind this was to make the process of finding the green beatnote easier. We used JAMMT (just another mode matching tool) to calculate a two lens solution to couple the light into the collimator - we use a +200mm and -200mm lens, I will upload a more detailed mode matching calculation + plot + picture soon. We wanted to have a beam waist of 350um at the collimator, a number calculated using the following formula from the Thorlabs website:

d =4\lambda (\frac{f}{\pi*MFD})

where d is the diameter of the output beam from the collimator, f is the collimating lens focal length and MFD is 6.6um for the fiber we use.

There is ~26mW of IR light coming through the BS after the EX AUX - after playing around with the 6 axis stage that the coupler is mounted on, Johannes got the IR transmission to the PSL table up to ~11.7mW. The mode matching efficiency of 45% is certainly not stellar, but we were more curious to find a beat and possibly measure the X arm loss so we decided to accept this for now - we could probably improve this by moving the lenses around. We then attenuated the input beam to the fiber by means of an ND filter such that the light incident on the coupler is now ~1.3mW, and the light arriving at the PSL table from the EX laser is ~550uW. Along with the PSL light, after the various couplers, we have ~500uW of light going to the IR beat PD - well below its 2mW threshold.

The IR beat was easily found with the frequency counter setup. However, there was no evidence of a green beat. So we went to the PSL table and did the near-field-far-field alignment onto the beat PD. After doing this, we were able to see a beat - but the amplitude was puny (~-60dBm, we are more used to seeing ~-20dBm on the network analyzer in the control room). Perhaps this can be improved by tweaking the alignment onto the PD while monitoring the RF output with an oscilloscope.

Moreover, the green PDH problems with the X end persist - even though the arm readily locks to a TEM00 mode, it frequently spontaneously drops lock. I twiddled around with the gain on the uPDH box while looking at the error signal while locked on a oscilloscope, but was unable to mitigate the situation. Perhaps the loop shape needs to be measured and that should tell us if the gain is too low or high. But ALS is getting closer to the nominal state...

Johannes is running his loss measurement script on the X arm - but this should be done by ~10pm tonight.

 

ELOG V3.1.3-